OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (unklerhaukus):

\[\large\quad-3+yi=x+6i\] lets compair the real parts \[~{\frak R}(-3+yi)={\frak R}(x+6i)\]\[\qquad\qquad-3=x\] and now the imaginary parts \[~{\frak I}(-3+yi)={\frak I}(x+6i)\]\[\qquad\qquad y=6\]

OpenStudy (anonymous):

that is one fancy \(\frak R\)

OpenStudy (unklerhaukus):